OpenStudy (majikdusty):

Mean value theorem question? determine whether f satisfies the hypotheses of the mean value on [a,b], and, if so, find all numbers c in (a,b) such that f(b) - f(a) = f'(c)(b-a). f(x)= tan(x) [0,(pi/4)] so i plugged the intervals in and got f(0)=0 and f(pi/2)=1 f'(c)= sec^2(c) Then i did (1-0)=sec^2(c)((pi/4)-0) ended up getting sec^2(c)=(4/pi) I'm not sure what to do from here... especially since a friend of mine got (4/pi) as an answer. Did i do something wrong? If not how do i finish this problem?

OpenStudy (anonymous):

i guess tangent is continuous on \([0\frac{\pi}{4}]\) so it does satisfy the hypothesis

OpenStudy (anonymous):

oh continuous and also differentiable

OpenStudy (anonymous):

also as you said \[f(\frac{\pi}{4})-f(0)=1-0=1\]

OpenStudy (anonymous):

ans also as you said the derivative is \(\sec^2(x)\) so your job is to solve \[\sec^2(x)=\frac{4}{\pi}\] good luck with that!